# The Preheating Stage on The Starobinsky Inflation after ACT

### v1

@@ 3 The preheating
where $\theta_{\mu}$ is a phase. It is shown that in Case 1, where the preheating is narrow, it should lie in the stability region, while Case 2 should belong to the instability region (see Fig. 2).
Figure 2: Mathieu instability region related to the semi-analytical result depicted in Eq. (3.14). The right legend corresponds to the value of $\mu_{p}$. Note that $Q$ can have negative values in our model. The white region corresponds to the stable condition.
4 The preheating duration and reheating
Generally, the preheating duration plays a role in determining the reheating temperature. If the duration is short, the reheating temperature is high, and vice versa [30]. In that case, a proper preheating duration is necessary for a viable reheating temperature. However, it is still model-dependent. Some peculiar models with long durations can also yield high reheating temperatures [^5] Higgs inflation could be one of the best examples. Due to the large non-minimal coupling, the duration of preheating should be long [30]. But its instant-type decay from gauge bosons to fermions gives a high reheating temperature, $T_{\text{reh}}\sim 10^{14}$ GeV [19].
To obtain the preheating duration, we will consider Eq. (2.3) and write the (oversimplified) e.o.m. together with the first and second Friedmann equations as [^6] The e.o.m. and two Friedmann equations here are rather oversimplified. As the purpose is to estimate the order of the time $t$, we think it is sufficient.
$\begin{split}&\ddot{\phi}+3H\dot{\phi}+M^{2}\phi=0,\\ &H^{2}=\frac{1}{3M_{p}^{2}}\left(\frac{1}{2}\dot{\phi}^{2}+\frac{1}{2}M^{2}\phi^{2}\right),\\ &\dot{H}=-\frac{\dot{\phi}^{2}}{2M_{p}^{2}}.\end{split}$ (4.1)
If we define $\dot{\phi}=\sqrt{6}HM_{p}\cos\theta$ and $M\phi=\sqrt{6}HM_{p}\sin\theta$, straightforwardly, we obtain
$\tilde{\phi}=2\sqrt{\frac{2}{3}}\frac{M_{p}}{Mt}.$ (4.2)
Lastly, following the method as Ref.[30], when the Jordan frame coincides with the Einstein frame, we may set $\tilde{\phi}_{\text{end}}=M_{p}$ and $\tilde{\phi}_{\text{reh}}=M$, yielding [31, 30]
$\Delta t=|t_{\text{reh}}-t_{\text{end}}|\simeq 2\sqrt{\frac{2}{3}}\frac{M_{p}}{M^{2}}\sim 10^{5}\pi M^{-1},$ (4.3)
where $t_{\text{reh}}$ is the time when the radiation-dominated era starts and $t_{\text{end}}$ is the time when preheating starts after the end of inflation. If we consider the result in the left figure of Fig. 1, the broad resonance still occurs at $t\gg 10^{5}\pi M^{-1}$. In that case, we conclude that such a mechanism is disfavored if preheating only adopts the Case 1-type. On the other hand, if the preheating only adopts Case 2, the preheating duration is too short compared to our analytical result in Eq. (4.3), so this is also disfavored. But realistically, during preheating, both $k_{\phi}$ and $k_{\chi}$ may not be fixed [30]. In this context, the dominant preheating process may arise from Case 1 being interrupted by Case 2 during the early or late stages of preheating. This combination allows for a preheating duration that aligns with the analytical result presented in Eq. (4.3). As will be discussed in Section 7, we will identify the specific time of the preheating period during which tachyonic preheating intervenes. [^7] In fact, we will get an indication that tachyonic preheating occurs during the initial stage of preheating from the discussion on PBH formation (see Section 6). To estimate the reheating temperature, it is necessary to write the formula as
where $a_{0}=1$ and $H_{0}=66.78\hskip 2.84526pt\text{km Mpc}^{-1}\text{s}^{-1}$ [13] represent the scale factor and Hubble parameter today. $a_{\kappa}$ and $H_{\kappa}$ represent the scale factor and Hubble parameter at the pivot scale $\kappa_{*}=0.05\hskip 2.84526pt\text{Mpc}^{-1}$. The subscripts ’eq’, ’end’, and ’reh’ correspond to ’matter-radiation equality’, ’end of inflation’, and ’reheating’, respectively. Note that we solve Eq. (4.4) by using $T_{\text{eq}}=0.8$ eV, and the end of preheating is evaluated at $\tilde{\phi}=M$. Finally, by using (4.4) with e-fold $N=75.5$, we obtain an upper bound for the reheating temperature as $T_{\text{reh}}\sim 10^{4}$ GeV. This value is five orders of magnitude lower than the previous prediction [2].
@@ 5 On the fermionic preheating
In this part, we will discuss the possibility that, instead of the scalar $\chi$, the scalaron is coupled with fermions. So far, we have only discussed the bosonic preheating, even though fermionic preheating is equally possible. This type of preheating has already been discussed in, e.g., Refs. [32, 33, 34, 35]. In this section, we will qualitatively discuss the possibility of this type of preheating.
If the scalaron is only coupled with certain fermions, at a large coupling [33], the preheating is similar to the bosonic case. However, it is heavily suppressed by the Pauli exclusion principle, in contrast to the Bose-Einstein condensate (BEC) that can occur during bosonic preheating. In this case, the inflaton energy transfer during resonance is mostly converted into the expansion of the universe. Fermion production becomes less suppressed once the universe expands. However, even after the end of preheating, the potential of the scalaron remains quadratic ($\propto\phi^{2}$), which still favors a matter-dominated regime. Fermions are more likely to be produced at the end of the preheating phase due to the looseness of the Pauli blocking.
Even if tachyonic preheating occurs, fermionic particle production remains heavily suppressed, unlike in the bosonic case. This means fermionic preheating is less efficient, even compared to Case 1 of bosonic preheating. Consequently, the duration of preheating is expected to exceed our analytical prediction. Finally, we safely said that preheating via this mode is strongly disfavored.
@@ 6 The PBH formation from preheating
[…]
@@
By using $\beta_{\text{PBH}}=10^{-20}$ and $\delta=\delta_{c}=0.7$ [36, 37] on Eq. (6.1), we obtain the mass variance $\sigma_{\delta}$ as
@@
For the production of PBH during preheating, the strong parametric resonance should be applied. Our assumptions implied that PBH is most likely produced when [38, 36]
$|Q|\equiv\left|\frac{2}{M^{2}}\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}-3k_{\phi}k_{\chi}\right)\frac{\tilde{\phi}}{M_{P}}\right|\gg p^{2}.$ (6.4)
The power spectrum of the variable $x$ can be expressed as [39]
$\mathcal{P}_{(x)}=\frac{k_{x}^{3}}{2\pi^{2}}\braket{}{x_{k}}{{}^{3}},$ (6.5)
where $k_{x}=|\textbf{k}_{x}|$ represents the comoving wavenumber associated with $x$. Note that $x_{k}$ denotes the inverse Fourier transform of $x$. The influence of preheating on $\chi$ production can be determined as [40, 38, 36]
$\mathcal{P}_{\delta\chi}=\mathcal{P}_{\delta\chi}|_{\text{end}}\exp\left(2\mu_{p}M(t-t_{\text{end}})\right).$ (6.6)
The characteristic exponent $\mu_{p}$ here is evaluated at $\theta_{\mu}=0$, in which $\mu_{p}=(2\pi)^{-1}\ln(1+2\exp(-\pi p^{2}))$, where $p$ is taken from Eq. (3.7) as
$p^{2}=\frac{k_{\chi}^{2}+m_{\chi}^{2}}{\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}-3k_{\phi}k_{\chi}\right)\frac{\tilde{\phi}}{M_{P}}M\right]^{2/3}}=\frac{1}{18\sqrt{Q}}\left(\frac{k_{\chi}^{2}+m_{\chi}^{2}}{k_{\text{end}}^{2}}\right),$ (6.7)
@@
$k_{\text{end}}^{2}=\frac{1}{18\sqrt{2}}\left|\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}-3k_{\phi}k_{\chi}\right)\frac{\tilde{\phi}}{M_{P}}\right]^{1/6}\right|M^{5/3}$ (6.8)
as the comoving wavenumber evaluated at the Hubble exit at the end of inflation. For $Q\gg 1$, in which it is required to get the largest production, we get $\mu_{p}=\mu_{0}\simeq\frac{1}{2\pi}\ln 3$ and our chosen parameters should be tachyonic ($m_{\chi}^{2}+3\xi M^{2}\ll 3k_{\phi}k_{\chi}$). In addition, we will calculate the non-adiabatic power spectrum of $\chi$ as [38, 36]
@@
[…]
Note that $\vartheta$ is the angle between $p$ and $p^{\prime}$. To solve the last equation, we need to impose [36]
@@
Finally, the mass variance $\sigma_{\delta}$ during the matter-dominated era can be calculated as follows [41]
@@
$5.43\times 10^{-3}\gtrsim\frac{e^{4\mu_{0}\cdot M\Delta t}}{(M\Delta t)^{3/2}}\cdot\frac{m^{4}_{\chi}}{M^{4}M_{p}^{2}}\frac{k^{\frac{5}{2}}_{\chi}}{\sqrt{k_{\phi}}}.$ (6.15)
Eq. (6.15) shows the constraints on $k_{\chi}$, $k_{\phi}$, $m_{\chi}$, and $\Delta t$ simultaneously. In the following, we describe the numerical results for Eq. (6.15) in Table 1. In this table, we set the parameters based on Case 2. The preheating duration is analytically found to be $\sim 10^{5}\pi M^{-1}$, and it is shown that a sufficient amount of PBHs is already produced at the early stage of preheating. However, due to the rapidly expanding universe, the overdense region is wiped out very quickly, leading to the rapid suppression of PBH formation as well [36].
$m_{\chi}$ $k_{\phi}$ $k_{\chi}$ $\Delta t$ ($\pi M^{-1}$)
$10^{-12}M_{p}$ $10^{-3.5}M_{p}$ $10^{-4}M_{p}$ $382$
$10^{-12}M_{p}$ $10^{-4}M_{p}$ $10^{-4}M_{p}$ $380$
$10^{-12}M_{p}$ $10^{-5}M_{p}$ $10^{-4}M_{p}$ $377$
$10^{-12}M_{p}$ $10^{-6}M_{p}$ $10^{-4}M_{p}$ $374$
$10^{-12}M_{p}$ $10^{-7}M_{p}$ $10^{-4}M_{p}$ $364$
$10^{-12}M_{p}$ $10^{-8}M_{p}$ $10^{-4}M_{p}$ $359$
Table 1: The duration of the required PBH abundance. The chosen parameters are assumed to follow the tachyonic conditions.
In our calculation of Eq. (6.15), the duration range to obtain a sufficient amount of PBH is always $\sim 10^{2}\pi M^{-1}$, insensitive to the chosen parameters of $m_{\chi}$, $k_{\phi}$, and $k_{\chi}$. In that case, the PBH constraint may not jeopardize our model.
7 The ceased-out mechanism
In this section, we will discuss the cease-out mechanism based on Ref. [30]. In this mechanism, the momenta $k_{\phi}$ and $k_{\chi}$ are not fixed; instead, they become time-dependent. As both momenta become functions of time, an appropriate functional form can be imposed to obtain the desired model.
In our previous calculation, it was shown that preheating in this model can be estimated to end at $\sim 10^{5}\pi M^{-1}$. Both cases indicate that neither mechanism alone can satisfy the prediction. Thus, the most plausible approach is to combine both mechanisms into one. In doing so, either $k_{\phi}$ or $k_{\chi}$ (or both) must be time-dependent.
In the numerical calculation of the equations (3.2) and (3.3), both $k_{\phi}$ and $k_{\chi}$ are fixed. However, in the cease-out mechanism, we assume $k_{\phi}$ to be a function of $t$ as
$k_{\phi}\equiv k^{0}_{\phi}e^{\alpha t},$ (7.1)
where $k^{0}_{\phi}$ is the initial amplitude of $k_{\phi}$ and $\alpha$ is a constant. As the production of $\chi$, especially during the first stage of preheating, should be heavy non-relativistic, it is implied that the momentum $k_{\chi}$ should be small.
In the left panel of Fig. 3, we show the condition when the ceased-out mechanism is imposed on $k_{\phi}$. In this numerical calculation, we set $\alpha=-10^{-5}$ in units of $M\pi^{-1}$. In this figure, $k_{\phi}$ is initially at $k_{\phi}^{0}=10^{-3.5}M_{p}$ but decreases exponentially. Consequently, the tachyonic growth stops after $t$ reaches $\sim 50000$ (in $\pi M^{-1}$). Thus, the time when preheating stops is favorable for our analytical prediction. On the other hand, in the right figure of Fig. 3, if we set $\alpha=10^{-5}M\pi^{-1}$ with initial momentum $k^{0}_{\phi}=10^{-5}M_{p}$, preheating is initially inefficient but becomes more efficient due to tachyonic preheating. At $t\sim 10^{5}$ (in $\pi M^{-1}$), preheating becomes narrow. However, to match the previous discussion about PBH, the left figure of Fig. 3 is disfavored. PBH formation should be significant in the early stage of preheating. Thus, a large momentum $k_{\phi}$ is necessary at the beginning of preheating.
Figure 3: The preheating of $\phi_{k}$ (black-dashed line) and $\chi_{k}$ (red-solid line) oscillations by momentum-dependent $k_{\phi}$ and $k_{\chi}$ affected by the ceased-out mechanism. In this numerical calculation of the left panel, we imposed $m_{\chi}=10^{-6}M_{p}$, $M=9.7\times 10^{-6}M_{p}$, $M_{p}=1$, $\xi=1$, $k_{\chi}=10^{-4}M_{p}$ and $k_{\phi}$ is written as Eq. (7.1) with $\alpha=-10^{-5}M\pi^{-1}$ and $k^{0}_{\phi}=10^{-3.5}M_{p}$. In the right panel, we used $k_{\phi}^{0}=10^{-5}M_{p}$ and $\alpha=10^{-5}M\pi^{-1}$; otherwise, all remained parameters are the same. In our numerical calculation, we also set $\phi_{k}(0)=M_{p}$, $\frac{d\phi_{k}}{dt}\big{|}_{t=0}=0$, $\chi_{k}(0)=0$, and $\frac{d\chi_{k}}{dt}\big{|}_{t=0}=0$.
@@ 8 Reheating scenario
In this section, we will discuss the possible reheating scenario in this model. It is noted that the reheating stage is the condition where the cosmological model can be closely linked to SM particle physics. However, it turns out that calculating the reheating temperature is not straightforward. We will see in the following subsection that several mechanisms can appear, and we will discuss which mechanism is the most favored.
@@ 8.2 The Perturbative decay of the scalaron to fermion $\psi$
By using the ACT reheating temperature constraint of $10^{4}$ GeV, $m_{\psi}\sim 10^{8}$ GeV could satisfy the prediction.
@@ 8.3 Remarks on the perturbative mode
In our discussion about the perturbative reheating by scalar $\chi$ and fermion $\psi$, it is shown that the scalar mode is disfavored by the ACT. Furthermore, as it is enhanced by BEC, the reheating temperature from this mode could be larger than $\sim 10^{9}$ GeV. On the other hand, the reheating mode via fermions is more favorable, as it can decay to SM particles due to their chiral nature [42].
In Ref. [43], it is shown that both reheating mechanisms are rather oversimplified, as they neglect the violation of the fluctuation-dissipation theorem (see Refs. [44, 45, 46]). Additionally, these mechanisms overlook the effects of preheating and can only be significant if preheating is inefficient. This regime slowly transitions to a radiation-dominated regime, and the perturbative decay channel opens only gradually. It is implied that perturbative reheating via bosons is strongly disfavored due to the high reheating temperature, which contradicts the low reheating temperature prediction. Also, during the matter-dominated regime, all dominant particle production is mostly heavy non-relativistic [40]. On the other hand, fermionic reheating suffers not only from the same problem but also from the Pauli exclusion principle, which further suppresses their production. Even though this mode is more favorable than the bosonic one, its reheating scenario is still not convincing.
In addition, Ref. [30], which discusses inflation with a non-minimal coupling model, shows that perturbative reheating is strongly disfavored. This is because achieving a favorable reheating temperature requires a violation of unitarity (i.e., $\xi\gg 1$).

### v3

@@ 4 The preheating
where $\theta_{\mu}$ is a phase. It is shown in the Mathieu instability band of Fig. 4 that certain parameters could potentially fall in the instability region in which the particle is largely produced. However, this semi-analytical result of the instability band could not show the effect of the initial conditions. Also, the parameter used in the numerical results could fit in the Fig. 4 with much smaller amplification by an order of 10.
Figure 4: Mathieu instability region related to the semi-analytical result depicted in Eq. (4.12). The right legend corresponds to the value of $\mu_{p}$. The white region corresponds to the stable condition.
@@ 5 On the fermionic preheating
In this part, we will discuss the possibility that, instead of the scalar $\chi$, the scalaron is coupled with fermions. So far, we have only discussed the bosonic preheating, even though fermionic preheating is equally possible. This type of preheating has already been discussed in, e.g., Refs. [greene2000fermionic, greene1999fermion, giudice1999fermion, peloso2000fermion]. In this section, we will qualitatively discuss the possibility of this type of preheating.
If the scalaron is only coupled with certain fermions, at a large coupling [greene1999fermion], the preheating is similar to the bosonic case. However, in our paper, it is heavily suppressed by the Pauli exclusion principle, in contrast to the Bose-Einstein condensate (BEC) that can occur during bosonic preheating. Fermion production becomes less suppressed once the universe expands. However, even after the end of preheating, the potential of the scalaron remains quadratic ($\propto\phi^{2}$), which still favors a matter-dominated regime. Fermions are more likely to be produced at the end of the preheating phase due to the looseness of the Pauli blocking. Finally, we safely said that preheating via this mode is strongly disfavored.
@@ 6 The PBH formation from preheating
$\beta_{\text{PBH}}=\frac{\rho_{\text{pbh}}}{\rho_{\text{tot}}}=\int^{\infty}_{\delta_{c}}P(\delta)d\delta,$ (6.1)
@@
By using $\beta_{\text{PBH}}=10^{-20}$ and $\delta=\delta_{c}=0.7$ [green2001primordial, niemeyer1999dynamics] on Eq. (6.1), we obtain the mass variance $\sigma_{\delta}$ as
@@
For the production of PBH during preheating, the strong parametric resonance should be applied. Our assumptions implied that PBH is most likely produced when [liddle2000super, green2001primordial]
$|Q|\equiv\left|\frac{2}{M^{2}}\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}\right)\frac{\tilde{\phi}}{M_{P}}\right|\gg A.$ (6.4)
The power spectrum of the variable $x$ can be expressed as [liddle2000cosmological]
$\mathcal{P}_{(x)}=\frac{k_{x}^{3}}{2\pi^{2}}\braket{|x_{k}|{}^{3}},$ (6.5)
where $k_{x}=|\textbf{k}_{x}|$ represents the comoving wavenumber associated with $x$. Note that $x_{k}$ denotes the inverse Fourier transform of $x$. The influence of preheating on $\chi$ production can be determined as [kofman1997towards, liddle2000super, green2001primordial]
$\mathcal{P}_{\delta\chi}=\mathcal{P}_{\delta\chi}|_{\text{end}}\exp\left(2\mu_{p}M\Delta t\right),$ (6.6)
where $\Delta t$ corresponds to the time required to obtain the appropriate number of PBH. The characteristic exponent $\mu_{p}$ here is evaluated at $\theta_{\mu}=0$, in which $\mu_{p}=(2\pi)^{-1}\ln(1+2\exp(-\pi p^{2}))$, where $p$ is taken from Eq. (4.7) as
$p^{2}=\frac{k_{\chi}^{2}+m_{\chi}^{2}}{\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}\right)\frac{\tilde{\phi}}{M_{P}}M\right]^{2/3}}=\frac{1}{18\sqrt{Q}}\left(\frac{k_{\chi}^{2}+m_{\chi}^{2}}{k_{\text{end}}^{2}}\right),$ (6.7)
@@
$k_{\text{end}}^{2}=\frac{1}{18\sqrt{2}}\left|\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}\right)\frac{\tilde{\phi}}{M_{P}}\right]^{1/6}\right|M^{5/3}$ (6.8)
as the comoving wavenumber evaluated at the Hubble exit at the end of inflation. For $Q\gg 1$, in which it is required to get the largest production, we get $\mu_{p}=\mu_{0}\simeq\frac{1}{2\pi}\ln 3$. In addition, we will calculate the non-adiabatic power spectrum of $\chi$ as [liddle2000super, green2001primordial]
@@
$I(p,M\Delta t)\equiv\frac{3}{2}\int^{\infty}_{0}p^{\prime 2}dp^{\prime}\int^{\pi}_{0}\sin\vartheta d\vartheta e^{2(\mu_{p^{\prime}}+\mu_{p-p^{\prime}})M\Delta t}.$ (6.11)
Note that $\vartheta$ is the angle between $p$ and $p^{\prime}$. To solve the last equation, we need to impose [green2001primordial]
@@
Finally, the mass variance $\sigma_{\delta}$ during the matter-dominated era can be calculated as follows [green2004NewCalculation]
@@
$0.0755^{2}\gtrsim\frac{0.02\cdot e^{0.6992\cdot M\Delta t}}{(M\Delta t)^{3/2}}\cdot\frac{m^{4}_{\chi}k_{\chi}^{3}}{\sqrt{\xi}M^{5}M_{p}^{2}}$ (6.15)
$m_{\chi}$ $k_{\chi}$ $\Delta t$ ($M_{p}^{-1}$)
$10^{-12}M_{p}$ $10^{-12}M_{p}$ $2.24\times 10^{7}$
$^{-12}M_{p} 1010$$^{-10}M_{p}$ $2.02\times 10^{7}$
$10^{-12}M_{p}$ $10^{-8}M_{p}$ $1.80\times 10^{7}$
$10^{-8}M_{p}$ $10^{-12}M_{p}$ $1.65\times 10^{7}$
$^{-8}M_{p} 1010$$^{-10}M_{p}$ $1.43\times 10^{7}$
$10^{-8}M_{p}$ $10^{-8}M_{p}$ $1.21\times 10^{7}$
$10^{-6}M_{p}$ $10^{-12}M_{p}$ $1.36\times 10^{7}$
$^{-6}M_{p} 1010$$^{-10}M_{p}$ $1.136\times 10^{7}$
$10^{-6}M_{p}$ $10^{-8}M_{p}$ $9.12\times 10^{6}$
Table 1: The duration of the required PBH abundance. In this table, we used $\xi=1$. Based on the table, it seems that the $\Delta t$ is insensitive to the chosen parameters.
Eq. (6.15) shows the dependence on $k_{\chi}$, $\xi$, $m_{\chi}$, and $\Delta t$. In the following, we present the numerical results for Eq. (6.15) in Table 1. From our numerical analysis, by varying $k_{\chi}$ and $m_{\chi}$ with fixed $\xi$, it appears that $\Delta t$ is insensitive to the chosen parameters. The duration required to obtain a sufficient amount of PBHs is found to be $\sim 10^{7}M_{p}^{-1}$. This value is consistent with Fig. 2, which shows that the effective growth of $\chi$ production occurs after $t>10^{6}M_{p}^{-1}$. Note that in our numerical calculation, we keep $m_{\chi},k_{\chi}\ll\xi M^{2}$. In this case, the PBH constraint may not jeopardize our model. Moreover, even if $\chi$ grows larger, at some point it cannot collapse into PBHs due to the rapidly expanding universe. Thus, PBHs are suppressed beyond a certain stage [green2001primordial].
@@ 7 Reheating scenario
In this section, we will discuss the possible reheating scenario in this model. It is noted that the reheating stage is the condition where the cosmological model can be closely linked to SM particle physics. However, it turns out that calculating the reheating temperature in 3 is not straightforward, but using the e-fold connection disregards the actual mechanism. We will see in the following subsection that several mechanisms can appear, and we will discuss which mechanism is the most favored.
@@ 7.2 The Perturbative decay of the scalaron to fermion $\psi$
By using the ACT reheating temperature constraint of $10^{4}$ GeV, $m_{\psi}\sim 10^{8}$ GeV could satisfy the prediction. Generally speaking, this mechanism is favored but we will see later that at some point, a certain mechanism is more favored.
@@ 7.3 Remarks on the perturbative mode
In our discussion about the perturbative reheating by scalar $\chi$ and fermion $\psi$, it is shown that the scalar mode is disfavored by ACT results as it produces a high reheating temperature. Furthermore, as it is enhanced by BEC, the reheating temperature from this mode could be larger than $10^{9}$ GeV. On the other hand, reheating via fermions is more favorable, as it naturally allows for lower temperatures and links the connection more into Standard Model particles due to their chiral nature [becker2024confronting].
In Ref. [allahverdi2010reheating], it is shown that both reheating mechanisms are rather oversimplified, as they neglect the violation of the fluctuation-dissipation theorem (see Refs. [hu1995fluctuation, berera2005absence, gleiser1994microphysical]). Additionally, these mechanisms overlook the effects of preheating and can only be significant if preheating is inefficient. This regime slowly transitions to a radiation-dominated regime, and the perturbative decay channel opens only gradually. It is implied that perturbative reheating via bosons is strongly disfavored due to the high reheating temperature, which contradicts our low reheating temperature prediction. Also, during the matter-dominated regime, all dominant particle production is mostly heavy non-relativistic [kofman1997towards]. On the other hand, fermionic reheating suffers not only from the same problem but also from the Pauli exclusion principle, which further suppresses their production. Even though this mode is more favorable than the bosonic one, its reheating scenario is still not convincing.
In addition, Ref. [risdianto2025second], which discusses inflation with a non-minimal coupling model, shows that perturbative reheating is strongly disfavored. This is because achieving a favorable reheating temperature requires violating unitarity.